# A Deep Automotive Radar Detector using the RaDelft Dataset

### v1

@@ VII Conclusions
This work introduces an innovative data-driven detector for automotive radar and the RaDelft dataset, a newly collected multi-sensor real-world dataset. The proposed detector is trained exclusively from unlabeled synchronized radar and lidar data, thus eliminating the need for costly manual object annotations. Two types of performance metrics were employed to validate the method, i.e., conventional probability of detection & probability of false alarm, alongside the Chamfer distance, a point cloud-level metric designed to capture spatial relationships and similarities between point clouds. The proposed method reduces by 5 meters (75% reduction) the Chamfer distance when compared with conventional OSCFAR detectors, and by 0.18 meters (10% reduction) when compared with the state-of-the-art. Also, it significantly increases the probability of detection. Moreover, an ablation study showed that including temporal information in the process is important, and Doppler information is especially crucial for our model’s good performance. Results show that the probability of detection is increased from 39.4% to 63.5%, and the Chamfer distance is reduced by 10% when using Doppler information.

### v3

@@ VII Conclusions
This work introduces an innovative data-driven detector for automotive radar and the RaDelft dataset, a newly collected multi-sensor real-world dataset. The proposed radar detector is trained exclusively from unlabeled synchronized radar and lidar data, thus eliminating the need for costly manual object annotations for the detection process. Two types of performance metrics were employed to validate the method, i.e., conventional probability of detection & probability of false alarm, alongside the Chamfer distance, a point cloud-level metric designed to capture spatial relationships and similarities between point clouds. The proposed method reduces by 4.2 meters (77% reduction) the Chamfer distance when compared with conventional OSCFAR detectors, and by 0.62 meters (28% reduction) when compared with the state-of-the-art. Also, it significantly increases the probability of detection. Moreover, an ablation study showed that including temporal information in the process is important, and Doppler information is especially crucial for our model’s good performance. Results show that the probability of detection is increased from 50.44% to 62.13%, and the Chamfer distance is reduced by 27% when using Doppler information.